# R u s s i a n   A c a d e m y   o f   S c i e n c e s  S i b e r i a n    B r a n c h

### v37

@@
Preface .  1)
Problems from the 1st Issue (1965) .  1)
Problems from the 2nd Issue (1966) .  1)
Problems from the 3rd Issue (1969) .  1)
Problems from the 4th Issue (1973) .  1)
Problems from the 5th Issue (1976) .  1)
Problems from the 6th Issue (1978) .  1)
Problems from the 7th Issue (1980) .  1)
Problems from the 8th Issue (1982) .  1)
Problems from the 9th Issue (1984) .  1)
Problems from the 10th Issue (1986) .  1)
Problems from the 11th Issue (1990) .  1)
Problems from the 12th Issue (1992) .  1)
Problems from the 13th Issue (1995) .  1)
Problems from the 14th Issue (1999) .  1)
Problems from the 15th Issue (2002) .  1)
Problems from the 16th Issue (2006) .  1)
Problems from the 17th Issue (2010) .  1)
Problems from the 18th Issue (2014) .  1)
Problems from the 19th Issue (2018) .  1)
New Problems (20th Issue, 2022).  1)
Archive of Solved Problems .  1)
Index of Names .  1)
@@ 5.36. What profinite groups satisfy the maximum condition for closed subgroups?
5.38. Is it the case that if $A,B$ are finitely generated soluble Hopfian groups then $A\times B$ is Hopfian? Peter M. Neumann
5.39. Prove that every countable group can act faithfully as a group of automorphisms of a finitely generated soluble group (of derived length at most 4). For background to this problem, in particular its relationship with problem 8.50, see (Peter M. Neumann, in: Groups–Korea, Pusan, 1988 ((Lect. Notes Math., 1398), Springer, Berlin, 1989, 124–139). Peter M. Neumann
@@
6.38. b) Is it true that an arbitrary subgroup of $GL_{n}(k)$ that intersects every conjugacy class is parabolic? How far is the same statement true for subgroups of other groups of Lie type? Peter M. Neumann
@@ 8.44. Prove or disprove that for all, but finitely many, primes $p$ the group
8.50. At a conference in Oberwolfach in 1979 I exhibited a finitely generated soluble group $G$ (of derived length 3) and a non-zero cyclic ${\mathbb{Z}}G$-module $V$ such that $V\cong V\oplus V$. Can this be done with $G$ metabelian? I conjecture that it cannot. For background of this problem and for details of the construction see (Peter M. Neumann, in: Groups–Korea, Pusan, 1988 ((Lect. Notes Math., 1398), Springer, Berlin, 1989, 124–139).
Peter M. Neumann
@@ 9.6. Is it true that an independent basis of quasiidentities of any finite group is finite?
9.32. What locally compact groups satisfy the following condition: the product of any two closed subgroups is also a closed subgroup? Abelian groups with this property were described in (Yu. N. Mukhin, Math. Notes, 8 (1970), 755–760).
Such groups are described (W. Herfort, K. H. Hofmann, F. G. Russo, Adv. Math., 390 (2021), 107894). Yu. N. Mukhin
@@
Comment of 2001: It is proved (S. Shelah, S. Thomas, Bull. London Math. Soc., 20, no. 4 (1988), 313–318) that the answer is positive in set theory with Martin’s Axiom. The question is still open in ZFC. Peter M. Neumann
9.40. Let $\Omega$ be a countably infinite set. Define a moiety of $\Omega$ to be a subset $\Sigma$ such that both $\Sigma$ and $\Omega\setminus\Sigma$ are infinite. Which permutation groups on $\Omega$ are transitive on moieties? Peter M. Neumann
@@
c) Does there exist a transitive permutation group on $\Omega$ that is transitive on $\aleph_{0}$-sections but which is a proper subgroup of ${\rm Sym}(\Omega)$? Peter M. Neumann
@@
9.42. Let $\Omega$ be a countable set and let $D$ be the set of total order relations on $\Omega$ for which $\Omega$ is order-isomorphic with $\mathbb{Q}$}. Does there exist a transitive proper subgroup $G$ of ${\rm Sym}(\Omega)$ which is transitive on $D$? Peter M. Neumann
@@ H. Zieschang
11.34. Describe the complex representations of quasisimple finite groups which remain irreducible after reduction modulo any prime number $q$. An important example: representations of degree $(p^{k}-1)/2$ of the symplectic group $Sp(2k,p)$ where $k\in{\mathbb{N}}$ and $p$ is an odd prime. Partial results see in (P. H. Tiep, A. E. Zalesski, Proc. London Math. Soc., 84 (2002), 439–472;  P. H. Tiep, A. E. Zalesski, Proc. Amer. Math. Soc., 130 (2002), 3177–3184). A. E. Zalesskiĭ
@@ B. Neumann
11.69. A group $G$ acting on a set $\Omega$ will be said to be 1-$\{2\}$-transitive if it acts transitively on the set $\Omega^{1,\{2\}}=\{(\alpha,\{\beta,\gamma\})\mid\alpha,\beta,\gamma\mbox{ distinct}\}.$ Thus $G$ is 1-$\{2\}$-transitive if and only if it is transitive and a stabilizer $G_{\alpha}$ is 2-homogeneous on $\Omega\setminus\{\alpha\}$. The problem is to classify all (infinite) permutation groups that are 1-$\{2\}$-transitive but not 3-transitive. Peter M. Neumann
@@
Peter M. Neumann, C. E. Praeger
@@
11.73. If a relatively free group is finitely presented, is it almost nilpotent?
@@ 11.84. Is the isomorphism problem soluble
(a) Yes, it is true (mod CFSG); moreover, if a locally finite group has such an involution, then it is finite of $n$-bounded order (S. V. Skresanov, https://arxiv.org/pdf/2402.05437.pdf).
(b) No, since $PSL_{2}(\mathbb{R})$ is an infinite simple group, which contains an involution but does not contain a subgroup isomorphic to the Klein 4-group (S. V. Skresanov, https://arxiv.org/pdf/2402.05437.pdf).
@@
[…]
@@
No, not always (P. Piwek, https://arxiv.org/abs/2403.02197).
@@ 12.53. Is it decidable whether or not two elements of a free $l$-group are conjugate?
Conjecture: No. Note that such a group cannot be 2-transitive (Gy. Károlyi, S. J. Kovács, P. P. Pálfy, Aequationes Math., 39 (1990), 161–166;  V. D. Mazurov, Siberian Math. J., 31 (1990), 615–617;  Peter M. Neumann, P. J. Rowley, in: Geometry and cohomology in group theory ((London Math. Soc. Lect. Note Ser., 252), Cambridge Univ. Press, 1998, 291–295). Peter M. Neumann
@@ A. Shalev
13.2. Does there exist a finitely based variety of groups $\mathfrak{V}$} such that the word problem is solvable in $F_{n}({\mathfrak{V}})$ for all natural $n$, but is unsolvable in $F_{\infty}({\mathfrak{V}})$ (with respect to a system of free generators)? M. I. Anokhin
@@
Yes, it is true (E. Khukhro, Preprint, 2025, http://arxiv.org/abs/2505.20999).
@@ 13.12. Is the group of all automorphisms of an arbitrary hyperbolic group finitely presented?
Comment of 2001: This is proved for the group of automorphisms of a torsion-free hyperbolic group that cannot be decomposed into a free product (Z. Sela, Geom. Funct. Anal., 7, no. 3 (1997) 561–593). O. V. Bogopol’skiĭ
@@
a) Yes, it is (A. Jaikin-Zapirain, Preprint, 2020, http://matematicas.uam.es/~andrei.jaikin/preprints/baumslag.pdf)
c) Yes, it is faithful (A. Jaikin-Zapirain, Preprint, 2020, http://matematicas.uam.es/~andrei.jaikin/preprints/baumslag.pdf).
@@
a) Yes, it exists (V. Arnautov, A. Topală, Bul. Acad. Ştiinţ. Repub. Moldova, Mat., 1998, no. 2(27), 130–131).
@@
d) No; moreover, no infinite abelian group satisfies this property with $k=2$ (Dekui Peng, Preprint, 2023, https://arxiv.org/abs/2310.08269).
@@ 14.21. Does a free pro-$p\thinspace$-group have the Noetherian Equation Property?
[…]
@@
A negative answer is equivalent to the fact that every linear pro-$p$-group satisfies a non-trivial pro-$p$-identity; this is known to be true in dimension $2$ (for $p\neq 2$: A. N. Zubkov, Siberian Math. J., 28, no. 5 (1987), 742–747;  for $p=2$: D. E.-C. Ben-Ezra, E. Zelmanov, Trans. Amer. Math. Soc., 374, no. 6 (2021), 4093–4128).
A. N. Zubkov, V. N. Remeslennikov
@@ 14.47. Is the lattice of all soluble Fitting classes of finite groups modular?
Comment of 2021: This is also proved for $n=3$ (A. Abdollahi, M. S. Malekan, Adv. Group Theory Appl., 13 (2022), 71–81). L. Levai, L. Pyber
@@ G. Fernández–Alcober
14.95. (C. R. Leedham-Green, Peter M. Neumann, J. Wiegold). For a finite $p\thinspace$-group $P$, denote by $c=c(P)$ its nilpotency class and by $b=b(P)$ its breadth, that is, $p^{b}$ is the maximum size of a conjugacy class in $P$. Class-Breadth Problem: Is it true that $c\leqslant b+1$ if $p\neq 2$?
@@
b) Yes, it is true for $n\geqslant 7$ (K. Kordek, D. Margalit, Bull. London Math. Soc., 54, no. 1 (2022), 95–111).
@@
Editors’ comment of 2021: an affirmative answer was obtained for the nilpotent case (N. Monod, J. Group Theory, 25, no. 5 (2022), 851–865). P. de la Harpe
@@ 15.13. Do there exist finitely presented branch groups?
Yes, there exist finitely generated non-amenable torsion branch groups (S. Kionke, E. Schesler, Bull. London Math. Soc., 56, no. 2 (2024), 536–550).

### v38

@@
Preface ........................................................................................................................................................................  1)
Problems from the 1st Issue (1965) ........................................................................................................................................................................  1)
Problems from the 2nd Issue (1966) ........................................................................................................................................................................  1)
Problems from the 3rd Issue (1969) ........................................................................................................................................................................  1)
Problems from the 4th Issue (1973) ........................................................................................................................................................................  1)
Problems from the 5th Issue (1976) ........................................................................................................................................................................  1)
Problems from the 6th Issue (1978) ........................................................................................................................................................................  1)
Problems from the 7th Issue (1980) ........................................................................................................................................................................  1)
Problems from the 8th Issue (1982) ........................................................................................................................................................................  1)
Problems from the 9th Issue (1984) ........................................................................................................................................................................  1)
Problems from the 10th Issue (1986) ........................................................................................................................................................................  1)
Problems from the 11th Issue (1990) ........................................................................................................................................................................  1)
Problems from the 12th Issue (1992) ........................................................................................................................................................................  1)
Problems from the 13th Issue (1995) ........................................................................................................................................................................  1)
Problems from the 14th Issue (1999) ........................................................................................................................................................................  1)
Problems from the 15th Issue (2002) ........................................................................................................................................................................  1)
Problems from the 16th Issue (2006) ........................................................................................................................................................................  1)
Problems from the 17th Issue (2010) ........................................................................................................................................................................  1)
Problems from the 18th Issue (2014) ........................................................................................................................................................................  1)
Problems from the 19th Issue (2018) ........................................................................................................................................................................  1)
New Problems (20th Issue, 2022)........................................................................................................................................................................  1)
Archive of Solved Problems ........................................................................................................................................................................  1)
Index of Names ........................................................................................................................................................................  1)
@@ 5.36. What profinite groups satisfy the maximum condition for closed subgroups?
5.38. Is it the case that if $A,B$ are finitely generated soluble Hopfian groups then $A\times B$ is Hopfian? P. M. Neumann
5.39. Prove that every countable group can act faithfully as a group of automorphisms of a finitely generated soluble group (of derived length at most 4). For background to this problem, in particular its relationship with problem 8.50, see (P. M. Neumann, in: Groups–Korea, Pusan, 1988 ((Lect. Notes Math., 1398), Springer, Berlin, 1989, 124–139). P. M. Neumann
@@
6.38. b) Is it true that an arbitrary subgroup of $GL_{n}(k)$ that intersects every conjugacy class is parabolic? How far is the same statement true for subgroups of other groups of Lie type? P. M. Neumann
@@ 8.44. Prove or disprove that for all, but finitely many, primes $p$ the group
8.50. At a conference in Oberwolfach in 1979 I exhibited a finitely generated soluble group $G$ (of derived length 3) and a non-zero cyclic ${\mathbb{Z}}G$-module $V$ such that $V\cong V\oplus V$. Can this be done with $G$ metabelian? I conjecture that it cannot. For background of this problem and for details of the construction see (P. M. Neumann, in: Groups–Korea, Pusan, 1988 ((Lect. Notes Math., 1398), Springer, Berlin, 1989, 124–139).
P. M. Neumann
@@ 9.6. Is it true that an independent basis of quasiidentities of any finite group is finite?
9.32. What locally compact groups satisfy the following condition: the product of any two closed subgroups is also a closed subgroup? Abelian groups with this property were described in (Yu. N. Mukhin, Math. Notes, 8 (1970), 755–760). Yu. N. Mukhin
Such groups are described (W. Herfort, K. H. Hofmann, F. G. Russo, Adv. Math., 390 (2021), 107894).
@@
Comment of 2001: It is proved (S. Shelah, S. Thomas, Bull. London Math. Soc., 20, no. 4 (1988), 313–318) that the answer is positive in set theory with Martin’s Axiom. The question is still open in ZFC. P. M. Neumann
9.40. Let $\Omega$ be a countably infinite set. Define a moiety of $\Omega$ to be a subset $\Sigma$ such that both $\Sigma$ and $\Omega\setminus\Sigma$ are infinite. Which permutation groups on $\Omega$ are transitive on moieties? P. M. Neumann
@@
c) Does there exist a transitive permutation group on $\Omega$ that is transitive on $\aleph_{0}$-sections but which is a proper subgroup of ${\rm Sym}(\Omega)$? P. M. Neumann
@@
9.42. Let $\Omega$ be a countable set and let $D$ be the set of total order relations on $\Omega$ for which $\Omega$ is order-isomorphic with $\mathbb{Q}$}. Does there exist a transitive proper subgroup $G$ of ${\rm Sym}(\Omega)$ which is transitive on $D$? P. M. Neumann
@@ 10.49. Does there exist a group $G$ satisfying the following four conditions:
Yes, it does exist (M. Amelio, Preprint, 2025, https://arxiv.org/abs/2509.11958).
@@ H. Zieschang
The project has been completed in (L. Di Martino, M. Pellegrini, A. Zalesski, J. Group Theory, 23 (2020), 235–285; A. Zalesski, European J. Math., 10 (2024), article no. 55; A. Zalesski, Preprint, 2025, https://arxiv.org/abs/2509.05770).
@@
11.34. Describe the complex representations of quasisimple finite groups which remain irreducible after reduction modulo any prime number $q$. An important example: representations of degree $(p^{k}-1)/2$ of the symplectic group $Sp(2k,p)$ where $k\in{\mathbb{N}}$ and $p$ is an odd prime. See partial results in (P. H. Tiep, A. E. Zalesski, Proc. London Math. Soc., 84 (2002), 439–472; P. H. Tiep, A. E. Zalesski, Proc. Amer. Math. Soc., 130 (2002), 3177–3184); C. Parker, M. van Beek, Preprint, 2024, https://arxiv.org/abs/2411.16379). A. E. Zalesskiĭ
@@ B. Neumann
Yes, there does (V. Bagayoko, https://arxiv.org/abs/2509.09186).
11.69. A group $G$ acting on a set $\Omega$ will be said to be 1-$\{2\}$-transitive if it acts transitively on the set $\Omega^{1,\{2\}}=\{(\alpha,\{\beta,\gamma\})\mid\alpha,\beta,\gamma\mbox{ distinct}\}.$ Thus $G$ is 1-$\{2\}$-transitive if and only if it is transitive and a stabilizer $G_{\alpha}$ is 2-homogeneous on $\Omega\setminus\{\alpha\}$. The problem is to classify all (infinite) permutation groups that are 1-$\{2\}$-transitive but not 3-transitive. P. M. Neumann
@@
Comments of 2025: The conjecture is proved in the case where $H$ acts innately transitively on the coset space $[H:U]$ (M. Fusari, A. Previtali, P. Spiga, J. Group Theory, 27 (2024), 929–965); for $n=3$ with $|H:U|\leqslant 10$ (L. Gogniat, P.  Spiga, Preprint, 2025, https://arxiv.org/abs/2502.01287); and if $H=UL$ where $L$ is a minimal $A$-invariant subgroup of $H$ (M. Fusari, S. Harper, P. Spiga, Bull. Austral. Math. Soc., 2025, https://doi.org/10.1017/S0004972725000176).
P. M. Neumann, C. E. Praeger
@@
11.73. If a relatively free group is finitely presented, is it virtually nilpotent?
@@ 11.84. Is the isomorphism problem soluble
(a) Yes, it is true (mod CFSG); moreover, if a locally finite group has such an involution, then it is finite of $n$-bounded order (S. V. Skresanov, J. Group Theory, 27, no. 6 (2024), 1197–1202).
(b) No, since $PSL_{2}(\mathbb{R})$ is an infinite simple group, which contains an involution but does not contain a subgroup isomorphic to the Klein 4-group (S. V. Skresanov, J. Group Theory, 27, no. 6 (2024), 1197–1202).
@@
    positive-definite if $\sum\limits_{k,\,l}f(g_{k}^{-1}g_{l})\,\overline{\!c}_{k}c_{l}\geqslant 0$ for any $g_{1},\ldots,g_{n}\in G$ and any
    $c_{1},\ldots,c_{n}\in{\mathbb{C}}$.
@@
No, not always (P. Piwek, Algebra Number Theory, 19, no. 8 (2025), 1663–1670).
@@ 12.53. Is it decidable whether or not two elements of a free $l$-group are conjugate?
Conjecture: No. Note that such a group cannot be 2-transitive (Gy. Károlyi, S. J. Kovács, P. P. Pálfy, Aequationes Math., 39 (1990), 161–166; V. D. Mazurov, Siberian Math. J., 31 (1990), 615–617; P. M. Neumann, P. J. Rowley, in: Geometry and cohomology in group theory ((London Math. Soc. Lect. Note Ser., 252), Cambridge Univ. Press, 1998, 291–295). P. M. Neumann
@@ A. Shalev
13.2. Does there exist a finitely based variety of groups $\mathfrak{V}$} such that the word problem is solvable in $F_{n}({\mathfrak{V}})$ for every positive integer $n$, but is unsolvable in $F_{\infty}({\mathfrak{V}})$ (with respect to a free generating system)? M. I. Anokhin
@@
a) Yes, it is true (E. Khukhro, to appear in Bull. London Math. Soc., 2025, http://arxiv.org/abs/2505.20999).
@@ 13.12. Is the group of all automorphisms of an arbitrary hyperbolic group finitely presented?
Comment of 2001: This is proved for the group of automorphisms of a torsion-free hyperbolic group that cannot be decomposed into a free product (Z. Sela, Geom. Funct. Anal., 7, no. 3 (1997) 561–593). O. V. Bogopolski
@@
a) Yes, it is (A. Jaikin-Zapirain, Ann. Sci. Éc. Norm. Supér. (4), 57, no. 4 (2024), 1101–1133).
c) Yes, it is faithful (A. Jaikin-Zapirain, Ann. Sci. Éc. Norm. Supér. (4), 57, no. 4 (2024), 1101–1133).
@@
The conjecture is proved (R. Kessar, G. Malle, J. London Math. Soc. (2), 111, no. 2 (2025), Article ID e70076, https://arxiv.org/abs/2311.13510).
@@
a) Yes, it is true (V. Arnautov, A. Topală, Bul. Acad. Ştiinţ. Repub. Moldova, Mat., 1998, no. 2(27), 130–131).
@@
b) The inequality does hold, but it is not sharp (V. I. Arnautov, Bull. Acad. Ştiinţe Repub. Moldova, Mat., 2010, no. 2 (2010), 3–19).
d) No; moreover, no infinite abelian group satisfies this property with $k=2$ (D. Peng, Preprint, 2023, https://arxiv.org/abs/2310.08269).
@@ 14.21. Does a free pro-$p\thinspace$-group have the Noetherian Equation Property?
14.37. Let $G(n)$ be one of the classical groups (special, orthogonal, or symplectic) of $(n\times n)$-matrices over an infinite field $K$ of non-zero characteristic, and $M(n)$ the space of all $(n\times n)$-matrices over $K$. The group $G(n)$ acts diagonally by conjugation on the space $M(n)^{m}=\underbrace{M(n)\oplus\cdots\oplus M(n)}_{m}$. Find generators of the algebra of invariants
$K[M(n)^{m}]^{G(n)}$.
@@
A negative answer is equivalent to the fact that every linear pro-$p$-group satisfies a non-trivial pro-$p$-identity. This is known to be true in dimension $2$ for $p\neq 2$ (A. N. Zubkov, Siberian Math. J., 28, no. 5 (1987), 742–747). It is also proved that a 2-dimensional linear pro-$2$-group in characteristic 2 satisfies a non-trivial pro-$2$-identity (D. E.-C. Ben-Ezra, E. Zelmanov, Trans. Amer. Math. Soc., 374, no. 6 (2021), 4093–4128). A. N. Zubkov, V. N. Remeslennikov
@@ 14.47. Is the lattice of all soluble Fitting classes of finite groups modular?
Comment of 2021: This is also proved for $n=3$ (A. Abdollahi, M. S. Malekan, Adv. Group Theory Appl., 13 (2022), 71–81). Comment of 2025: Significant progress was obtained in (S. Kionke, N. Otmen, T. Toti, M. Vannacci, T. Weigel, Preprint, 2025, https://arxiv.org/pdf/2507.19086). L. Levai, L. Pyber
@@
a) These numbers are found (R. I. Gvozdev, Ya. N. Nuzhin, Siberian MAth. J., 64, no. 6 (2023), 1160–1171; M. A. Vsemirnov, R. I. Gvozdev, Ya. N. Nuzhin, Abstracts of Int. Conf. Mal’cev Meeting 2023, Novosibirsk, 2023, p. 148).
@@ G. Fernández–Alcober
14.95. (C. R. Leedham-Green, P. M. Neumann, J. Wiegold). For a finite $p\thinspace$-group $P$, denote by $c=c(P)$ its nilpotency class and by $b=b(P)$ its breadth, that is, $p^{b}$ is the maximum size of a conjugacy class in $P$. Class-Breadth Problem: Is it true that $c\leqslant b+1$ if $p\neq 2$?
@@
b) Yes, it is true for $n\geqslant 5$ (S. Orevkov, Ann. Fac. Sci. Toulouse. Math., 33, no. 1 (2024), 105–121, extending the result for $n\geqslant 7$ in K. Kordek, D. Margalit, Bull. London Math. Soc., 54, no. 1 (2022), 95–111).
@@
Comment of 2021: an affirmative answer was obtained for the nilpotent case (N. Monod, J. Group Theory, 25, no. 5 (2022), 851–865). Comment of 2025: this problem is reduced to the case of simple Lie groups; in particular, the solvable case is settled (A. Conversano, N. Monod, J. Algebra, 640 (2024), 106–116). P. de la Harpe
@@ 15.13. Do there exist finitely presented branch groups?
b) Yes, there exist finitely generated non-amenable torsion branch groups (S. Kionke, E. Schesler, Bull. London Math. Soc., 56, no. 2 (2024), 536–550).